Session story · 2026-09-05-a

The sweep offset was ten flag bits

Session: 2026-09-05, 04:29 → 08:30 UTC (two Claude Code conversations, one bench session; the second conversation opened while Chris was running the procedure the first one wrote) · Commits: wip/octave-flags-20260905 → d5ffdaa (parent), 5f75f9e (Cascade)

Where we started

FFT decode had been validated against APWIN. Claude's queue had five items, among them a cable A/B, generator level snapping in cal gaps, and a Ch B decode bug. The app had no CSV export, so two runs could only be compared by eye.

The working belief about the cables was that a 12 ft run could not change the level, and that any difference would be hum picked up where the RCA-BNC adapters break the balanced path.

Chris opened with: "ok - back at it after a while. what is next? By the way, I have a straight through XLR cable now. as part of our work I would like to test it against the longer ones with adapters. This is a 4 foot cable. the setup we have been using is 2 that are about 6 feet each, joined with RCA-BNC adapters"

The question

Does the adapter chain change what the analyzer measures? Underneath it, a second question appeared during the session: why does the DUT-mode frequency sweep read 4.61 dB higher than the FFT?

What happened

Tools first. Chris chose CSV export first and the Ch B fix afterwards. Claude built Cascade/CSVExport.swift, docs/compare_cables.py and an A→B→A procedure. It also fixed two Ch B bugs by reasoning alone, said neither had touched hardware, and added a CH B RANGING probe.

The first results, and a script bug. The frequency-response comparison printed "no overlapping frequencies" for every run. Claude's interp() assumed ascending frequencies, and the sweep ran 20 kHz → 20 Hz. With that fixed, the level answer was clean: B − A1 was −0.006 dB on the FFT fundamental and +0.006 dB on the sweep mean, against a control (A2 − A1) of −0.013 and −0.021 dB. Claude also reported the adapter chain 5–7 dB worse on hum, using integrated bands because the A-vs-A control had swung 16.8 dB on one bin. And the sweep read +4.47 dBV where the FFT read −0.13, a 4.61 dB gap. Claude attributed that to a missing DUT correction and declined to fix it from one operating point.

The garbage FFT. Chris had left fft-first-garbage.csv among the exports. Claude added a detector and wrote that the cause was unknown, reasoning from the CSV timestamps that a clean capture had also been a first-FFT-after-a-sweep. Chris corrected the premise: "the "garbage" one I was going to give you in a different question. That happened twice, right after the packats get loaded in, on the first run of FFT after it prepped itself." With that, Claude found the mechanism in the code. The T69 blob leaves the analyzer on a code-3 range sized for −20 dBV, and the acquisition slept a fixed 0.4 s while autorange climbed out of a 16 dB overload. The fix polls until the range settles. Claude's summary of its mistake: the files someone keeps are not a log of what they did.

The sweep direction. A queued message from Chris: "when in the interface I say "start frequency 20Hz, stop frequency 20kHz", it still runs from start 20kHz and stop 20Hz. I think this was something you fixed earlier. maybe it regressed or there's some other issue." It had never been wired: since 2026-08-23 the fields set span and point count, never direction.

The retraction. "OK - i took one FFT test right after the blob load and it looks good. CSV at fft-retest" The retest was clean, but its hum matched config B's to within ±0.37 dB. Claude asked which cable had been on it; Chris answered the 4 ft straight XLR. Claude retracted the hum finding. A→B→A had produced state 1 → 2 → 1, which tracked the cable only by order. Claude wrote the rule down: interleave A/B/A/B in one session on one build. The level result stood, four fundamentals within 0.018 dB.

Chris's hypothesis. "yous aid " The sweep's DUT-mode level reads 4.61 dB high" --that test measured both channel A and Channel B. Channel B was at 0 in the graph but there was no cable attached at all. Is it possible that 4+dB problem with Channel A was due to some reference against the unplugged chb?" Claude found a plausible route through the shared octave register EB 02 02, flagged that the simplest version predicted the wrong sign, added a console check, and wrote an interleaved B-open / B-connected procedure.

Making the procedure runnable. Chris asked for concrete steps: "so to be clear, what exact procedure do you want run now?", then "i don't really know what position that pot needs to be in to get -15dB". Claude made the probe measure the attenuation itself. Asked "is bench_chb_and_sweep_offset.md accurate?", Claude checked it against the code and found one error: the sweep's cal always jumps the generator to 0.0 dBV. Then "are you able to write this console log to a file yourself?" and "are you able to automate the csv exports?" produced a stdout tee and a session folder of auto-exported CSVs with a RUN LABEL field.

The run. In the second conversation Chris reported the tests done and asked about the CH B RANGING output: "tell me if i am wrong though". It was in console.log, lines 510–541, because the tee captures all of stdout. The probe showed the rangers are independent, so the Ch B exponent fix was aimed at a real bug. It also showed that oct = codeA × 1024 + 62 on all 13 rows.

The cause. EB 02 02 took three values all session: 6158 (B connected), 6918 and 6662 (B open). All three are range code 6 when shifted right by 10. The differences were in the low 10 bits, which were being decoded as magnitude at 0.0058841 dB/unit. Prediction against measurement: +4.472 vs +4.472 dB, +2.966 vs +2.966 dB, and a bimodal s3 sweep with 28 points at 4.474 and 2 at 2.966. Ch B open was a trigger that changed the flags. It was not the cause. Chris added: *"i have to note here

absolute level with no external check anywhere in the session.

Side issues. The probe's log had pushed the sidebar navigation off the window; Claude bounded it. Chris noticed "the default Run Label is "s1_Bopen" ." It was a placeholder Claude had written, so stage 1 exported unlabeled. Claude showed stage 1 was still identifiable from the log, and exports now write (unset).

Confirmation, on the second try. Chris first ran an FFT by mistake. Its cal line showed reg02=6144, exactly 6 × 1024, so the mask was live. Claude said it was not yet the confirmation run. "right. i mistakenly ran FFT instead of frequency response." The sweep followed: −0.092 … 0.000 dBV with Ch B open, octA 6144 on all 33 reads, and the octave warning silent. The same FFT log showed a headline THD+N floor of −57.1 dB against a robust floor of −104.5 dB. The integral started at 0 Hz while the bands started at 20 Hz. Claude fixed it, predicting −87.1 dB and noting the remaining gap is window sidelobes.

A second correction by Claude. Claude re-checked the probe's −30 row it had called a cal-gap snap; it was the ranger stale after a 30 dB step. The real cal-gap problem was wider than stated: seven cell boundaries, 6 of 37 points mis-plotted at 1 dB steps. Claude added a guard that leaves a hole instead of a wrong point, and wrote T70 and T71 procedures once Chris asked whether one existed.

Ending. "ok i am going to restart the machine, because stopping the cascade app has hung. write this to memory so we can resume after restart" Claude wrote the resume memory and said nothing was committed. "should you commit now before restart?" Claude found the submodule in detached HEAD, branched both repos and committed there.

The turns

  1. "That happened twice, right after the packats get loaded in" replaced a reconstruction from file timestamps with what actually happened.
  2. Chris named the cable on the retest, which broke the A→B→A correlation.
  3. "Is it possible that 4+dB problem with Channel A was due to some reference against the unplugged chb?" The hypothesis was wrong, but the experiment it prompted found the cause.
  4. Interleaving: the B-open runs disagreed with each other, which is what exposed flag bits.

Beliefs that changed this session

BelievedThenWhat overturned it
the adapter chain adds 5–7 dB of humtwo states, not tied to the cableretest on cable A matched B's hum
garbage FFT cause unknownfirst FFT after blob clips on code 3Chris: it happened twice, first run after load
sweep direction regressedit was never wiredcode history from 2026-08-23
4.61 dB is a missing DUT correctionlow 10 bits of EB 02 02 are flagsthree register values, exact dB prediction
open Ch B shifts the shared octaveCh B open only changes the flagsoct >> 10 == codeA on all 13 probe rows
the −30 probe row was a cal-gap snapthe ranger was stalethe next row at −27 read −26.64
cal gaps: below −30 and 4–6 dBVseven boundariesre-reading the table cells
RUN LABEL defaults to s1_Bopenit was a placeholderChris, then AppState.swift:25

What we learned about the instrument

What we learned about working together

Evidence

Where we ended

The sweep offset is fixed and confirmed on hardware. The first-FFT clipping fix held on five captures. Work sat on wip/octave-flags-20260905, not on main. Open: run T70; the stereo FFT capture (Ch2Source = 1), which would give Ch B an external check; the hum states; the FFT blob reload time Chris raised ("the APWIN program is quite fast in doing FFTs"); and the stop hang that ended the session.